1. Two persons A and B purchased 2kg of rice and 3kg of wheat for Rs. 86, 4 kg of rice and 5kg of wheat for Rs. 140. Find the price of 1 kg of wheat.

Answer by edjones(7311)

A) 2r+3w=86
B) 4r+5w=140
.
A)-4r-6w=-172 multiply each side by -2
B) 4r+5w=140
_____________Add
=====>-w=-32
.
w=Rs. 32
.
Ed